OpenStudy (anonymous):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x 8 (Points : 4) parallel perpendicular neither parallel nor perpendicular they are the same line

OpenStudy (anonymous):

6y=-3x+8 y=-3/6+8/6 y=-1/2+8/6 since the slop is reciprocal, it is perpendicular